OpenStudy (anonymous):

the floor area of a home is 1827 sq ft. Find the area in square meters. We are going over conversion factors and I am having trouble setting up the "road map" in my chemistry problem. Im confused what to square what not to square within the conversion factor

OpenStudy (neer2890):

1feet= 0.30480 m

OpenStudy (neer2890):

1827 sq ft= (1827*0.30480) sq. m

OpenStudy (anonymous):

Hi I was wondering if you know how to set up the conversion factor for the problem. The conversion factor is basically the fraction 1827 is multiplies by in order to get the answer

OpenStudy (neer2890):

sq. feet or sq. m only represents that it is area of a square

OpenStudy (anonymous):

wold you happen to know if I have to go from feet to centimeters to meters. I'm suppose to convert using the metric system

OpenStudy (neer2890):

ohk.. in that case 1m= 3feet and 27/8 inches

OpenStudy (jtvatsim):

Here's how I would approach it: \[\frac{1827 \ \text{ft}^2}{1} \cdot \frac{0.3048 \ \text{m}}{1 \ \text{ft}} \cdot \frac{0.3048 \ \text{m}}{1 \ \text{ft}}\] You need two multiplications by 0.3048 in order to get meters squared.

OpenStudy (anonymous):

ok I don't have to include centimeters?

OpenStudy (jtvatsim):

You would only need to include centimeters for two reasons: 1) The answer asked for square centimeters. 2) You only have the conversion fact from feet to centimeters or inches to centimeters. Some chemistry courses only tell you that 1 in = 2.54 cm. If that is all you have to work with you have to convert feet to inches, inches to centimeters, and then centimeters to meters. However, we have given you a new conversion fact 1 foot = 0.3048 meters which allows you to take a shortcut. Does that make sense?

OpenStudy (anonymous):

ok yes, basically use the information given. Quick question tho if I had used centimeters which I know 1meter=10^2cm would there be a need to square the centimeter? Its just for my own future reference

OpenStudy (jtvatsim):

Yes, you would need to square the centimeter. I'll type the reason below.

OpenStudy (jtvatsim):

\[\frac{1 \ m^2}{1} \cdot \frac{100 \ cm}{1 \ m} \cdot \frac{100 \ cm}{1 \ m} = 10000 \ cm^2\]

OpenStudy (jtvatsim):

I guess I should have asked if you were familiar with this kind of conversion approach? Otherwise that explanation might not have made sense. :)

OpenStudy (anonymous):

no it does make sense that is how they want us to do it but when I was doing it on my own I was going from ft sqrd to cm(which I didn't know to square or not) and then since I knew 10^2cm= 1meter

OpenStudy (anonymous):

Its that middle part to square the cm or not that was making me confused (if I do it in 3 steps vs 2)

OpenStudy (jtvatsim):

I see. Basically, anytime you see a squared, it means you probably have to use your conversion fact twice.

OpenStudy (jtvatsim):

You could also use this method if you prefer.

OpenStudy (jtvatsim):

1 m = 100 cm (1 m )^2 = (100 cm)^2 1^2 m^2 = 100^2 cm^2 1 m^2 = 100^2 cm^2

OpenStudy (jtvatsim):

You basically just square both sides to create a new conversion fact for squares.

OpenStudy (jtvatsim):

It works just as well. But use whatever makes sense to you and that will make your teachers happy. :)

OpenStudy (anonymous):

Thank you so much! :) so square everything in between even if I'm doing it in 3 steps! Thanks :)

OpenStudy (jtvatsim):

Yes, assuming that you are dealing with squares at the beginning and at the end. :)